OpenStudy (anonymous):

solve for x :| 7-x|+2=12

OpenStudy (anonymous):

Start with moving 2 to the right side:

OpenStudy (anonymous):

x = -3 |7-x|+2 = 12 |7-x| = 10 7-x = 10 -x = 3 x = -3 X = -3 but the process of removing the absolute value may be wrong. its an assumption.

OpenStudy (anonymous):

@franciscanmonk another case of absolute value 7-x = -10 ...

OpenStudy (anonymous):

ah haha. i knew i forgot something.